A bullet moves in a circular path and its rotational velocity w = πt + 3 / 2π after t = 0, how many seconds does it take for the bullet to complete one round?


1
Expert's answer
2021-02-04T19:09:16-0500

With given rotational velocity the covered angle is:


"\varphi(t) = \int_0^t\omega(t)dt = \int_0^t\pi t+1.5\pi dt =\dfrac{\pi t^2}{2} + 1.5\pi t"

Here we assumed that "\varphi(0) = 0". After one revolution, the agle becomes "\varphi(t) = 2\pi" substituting it into the formula above obtain the equation for "t":


"\dfrac{\pi t^2}{2} + 1.5\pi t = 2\pi\\
0.5t^2+1.5t-2 = 0\\
t^2 +3t-4 = 0"

The roots are:


"t_1 = 1, t_2 = -4"

Taking the positive root (since we consider time after t=0), find the final answer.


Answer. 1s.